M and N are the midpoints of the diagonals AC and BD respectively of quadrilateralABCD, then AB+ AD+CB+CD = _______________

Grade:12

1 Answers

Ujjawal maru
6 years ago
In a quadrilateral ABCD given that M is the midpoint of a vector and C vector which is equal to a vector + c vector divided by 2 and similarly N is equal to b vector + d vector divided by 2 now now considered a vector is a ,b vector as b C vector as c and d vector as D. A B + AD + CB +CD = b-a+d-a+b-c+d-c=2[(b+d)-(a+c)]=4(N-M)